# Stochastic Dominance Constrained Optimization with S-shaped Utilities: Poor-Performance-Region Algorithm and Neural Network

### v1

@@ 4.2 SSD Problem: Numerical Algorithm
    In Steps 8-12, we need to check whether the constructed $y_{\text{sub}}(\cdot)$ is non-decreasing on (0,1) to make the solution valid. It holds in many cases. However, in some extreme cases, we could still ensure the monotonicity by redesigning the construction of $t_{i}$ defined in Eq. (20), which is named $t_{\text{left}}$ in Eq. (23). The specific case will be shown in Section 5.
@@
    In Step 15, we need to check the condition holds numerically. It holds in many cases.
@@ 5 Numerical Results
Our study is motivated by the Black-Scholes model in a complete market. A classic Black-Scholes model consists of one riskless bond ($\frac{\mathrm{d}B_{t}}{B_{t}}=r\mathrm{d}t,~t\in[0,T)$, where the risk-free rate is $r=0.05$) and one stock ($\frac{\mathrm{d}S_{t}}{S_{t}}=\mu_{\text{S}}\mathrm{d}t+\sigma_{\text{S}}\mathrm{d}W_{t},~t\in[0,T)$, which is a geometric Brownian motion with the expected return rate $\mu_{\text{S}}=0.086$ and the volatility parameter $\sigma_{\text{S}}=0.3$ and $\{W_{t}\}_{0\leqslant t\leqslant T}$ is a standard Brownian motion). The wealth process is given by $\mathrm{d}X_{t}=\left(rX_{t}+(\mu_{S}-r)\pi_{t}\right)\mathrm{d}t+\sigma_{S}\pi_{t}\mathrm{d}W_{t},~t\in[0,T)$ and $X_{0}=\overline{x}$, where $\{\pi_{t}\}_{0\leqslant t<T}$ is the control process representing the investment amount in the stock and $T\in(0,\infty)$ is the evaluation time of investment. We define the pricing kernel process $\{\rho_{t}\}_{0\leqslant t\leqslant T}$ by
@@
As the market is complete and one can use the martingale method to duplicate the optimal portfolio process, it is sufficient to solve the optimal terminal wealth variable via the static problem (1) (see, e.g., Liang and Liu (2024)). Hence, our focus is solving the optimal wealth variable in Problem (1). Adapting to Problem (1), we denote the terminal variables of pricing kernel and wealth by $\rho:=\rho_{T}$ and $X:=X_{T}$, with a slight abuse of notation. Solving from Eq. (29), $\rho$ follows the log-normal distribution (i.e., $\log(\rho)\sim\text{N}(-(r+\theta^{2}/2)T,(\theta\sqrt{T})^{2})$) and has a quantile function
$Q_{\rho}(t)=\exp\left\{\theta\sqrt{T}\Phi^{-1}(t)-(r+\theta^{2}/2)T)\right\}\triangleq\exp\left\{\sigma\Phi^{-1}(t)+\mu\right\},~t\in(0,1),$ (30)
@@ 5.1 FSD Problem: S-shaped Utility (Theorem 1)
We begin by specifying a S-shaped utility $U:[L,\infty)\rightarrow\mathbb{R}$, following the general S-shaped utility formulation in Definition 2:
@@
We then consider two portfolio selection problems. The first is Problem (3) with the general setting of S-shaped utility in Definition 2 (In particular, Eq. (31)) and the benchmark quantile $Q_{0}(t)=10t^{2}-1,\;t\in[0,1].$ The second is the Merton problem (1) using the S-shaped utility setting in Eq. (31):
[…]
where the liquidation boundary is given by L=5L=-5. Here we add the liquidation boundary in order to make the second problem well defined and compare with the first problem.
In this example, the first-order SD constraint acts a similar role as the liquidation constraint: if $Q_{0}(1-F_{\rho}(\rho))=Q_{0}(0)$, then the optimal solution $X^{*}$ locates at the boundary $Q_{0}(1-F_{\rho}(\rho))$, otherwise $X^{*}$ is the same as the classic solution $(U_{1}^{\prime})^{-1}(\lambda\rho)$.
Figure 1: First-order SD constraint: $\bar{x}=5$
@@ 5.2 SSD Problem: Power Utility
$Q_{0}(t)=\exp\left\{\sigma_{0}\Phi^{-1}(t)+\mu_{0}\right\},\;t\in[0,1].$
@@
Table 1: Parameter setting in the numerical illustration.
@@
Figure 2: Impacts of $\mu_{0}$ and $\sigma_{0}$.
@@
We apply the Poor-Performance-Region Algorithm (PPRA), the explanations and financial insights from Figure 2 and Table 2 are as follows:
@@ 5.3 SSD Problem: S-shaped Utility
Table 3: Parameter setting in Figure 3.
@@
We apply the Poor-Performance-Region Algorithm (PPRA), the explanations and financial insights from Figure 3 and Table 4 are summarized as follows:
@@
    Figure 3: Across the S-shaped utility cases, we observe that when the market performs poorly, relative to the classical solution, the PPRA wealth exhibits a clear improvement, driven by the SSD constraint. Therefore, in adverse market scenario, the SSD constraint effectively performs as a risk-control mechanism, ensuring that the PPRA wealth remains at least as high as the benchmark wealth.
@@
    Figure 3 (e) vs (f): In these examples, the poor performance region splits into two disjoint intervals. We further observe that when the budget of $Q_{0}$ is close to the bound $\bar{x}$, the PPRA wealth becomes increasingly close to the benchmark. This phenomenon arises because a high level of the budget of $Q_{0}$ make the SSD constraint dominate the optimization.
@@ 5.4 SSD Problem: Various Utilities and Benchmark Quantiles
Based on the proposed algorithm, we further extend its applicability to a broader class of the SSD problem. To assess the generality of our approach, we conduct numerical experiments using different utilities and benchmark quantiles $Q_{0}$.
For the utility function $U(x)$, we consider several representative forms capturing different risk preference, as shown in Table 5.
@@
Table 5: Setting of various utilities.
@@
Table 7: General setting in Figures 4-5.
@@
(e) piecewise log-normal 3.5 $\mu_{0}=-1,\sigma_{0}=3,k_{0}=2.3$
$p_{1}=q_{1}=0.6,p_{2}=0.8,$
$q_{2}=0.9,\lambda_{1}=1,\lambda_{2}=2$
(f) piecewise exponential 1.3 $\alpha=0.7,k_{0}=2.3$
$p_{1}=q_{1}=0.6,p_{2}=0.8,$
$q_{2}=0.9,\lambda_{1}=1,\lambda_{2}=2$
Table 8: Parameter setting in Figures 4-5.
Applying the Algorithm 1, we obtain the numerical results in Table 9 and plot Figures 4-5, where each panel illustrates the structure of the correction function $y_{\text{sub}}(\cdot)$ and the Poor-Performance-Region Algorithm (PPRA) solution.
@@
Figure 5: Setting (d)-(e)-(f).
Later, we would clarify the structural behavior of the correction function $y_{\text{sub}}(\cdot)$. The numerical experiments yield the following explanations and financial insights:
@@
    Figure 4 (a)-(b)-(c): In bad market scenario, the correction function $y_{\text{sub}}(\cdot)$ satisfies $y_{\text{sub}}(\cdot)=y_{0}(\cdot)$, therefore, the PPRA wealth will coincide with the benchmark wealth. In this scenario, the benchmark serves as a risk-control mechanism, which is similar to Figure 2 (a)-(b)-(d).
@@
    Figure 4 (c)-(f): Compared with Figure 4 (a)-(b), the region where $y_{\text{sub}}(\cdot)=y_{0}(\cdot)$ expands. Hence, under relatively stagnant or bad market scenario, the PPRA wealth tends to coincide with the benchmark wealth. In such scenario, the SSD constraint plays a more prominent role by ensuring a benchmark wealth to reduce risk.
@@
    Figure 5 (d): The correction function $y_{\text{sub}}(\cdot)=y_{0}(\cdot)$ over certain intermediate intervals, causing the PPRA wealth to coincide with the benchmark in mid-range market scenario, as shown in (d)-Right.
@@
    Figure 5 (e): The correction function $y_{\text{sub}}(\cdot)=y_{0}(\cdot)$ at both endpoints of the interval $(0,1)$. This implies the PPRA wealth must coincide with the benchmark wealth in both extremely bad and exceptionally favorable market scenarios.
In Figure 5 (e), we can observe how Steps 8-12 of Algorithm 1 ensure that the constructed function $y_{\text{sub}}(\cdot)$ remains non-decreasing on $(0,1)$. The key idea is to modify the construction of $g_{i}(t)$ in Eq. (19) so that both monotonicity requirement of $y_{\text{sub}}(\cdot)$ and the condition $z^{*}(\cdot)\leqslant 0$ is satisfied throughout the region in which $y_{\text{sub}}(\cdot)$ is updated. In this example, the interval where $y_{\text{sub}}(\cdot)=y_{\text{sub}}(t_{1})$ is extended into the region covered by the previous iteration (i.e., the first iteration), updating the original $y_{\text{sub}}(\cdot)$ in this region.
@@ 6 Algorithm-Guided Piecewise-Neural-Network Framework
In this section, we propose a novel approach to solve the SSD Problem (14) by designing a piecewise-neural-network-framework combined with Algorithm 1. The key observation is that in Algorithm 1, the poor performance region $C$ defined in Eq. (27) and the construction of the correction function $y_{\text{sub}}(\cdot)$ provide valuable structural information about the optimal solution, thereby guiding the design of an effective neural network framework.
@@ 6.1 Model Setting
where $f^{(k)}_{\theta}$ denotes the $k$-th neural sub-network parameterized by $\theta_{k}$, the parameters of the integrated network $Q$ are denoted as $\theta=\{\theta_{0},\theta_{1},\dots,\theta_{K}\}$. $\mathbf{s}_{\text{feat}}(s)$ represents the Fourier feature embedding of the scalar input $s$ defined as
@@
$C_{1}=\frac{1}{n}\sum_{i=1}^{n}Q_{\theta}(s_{i})Q_{\rho}(1-s_{i}),\;\;C_{2}=\max\left\{0,\max_{k=1,\dots,n}\left[\frac{1}{n}\sum_{i=1}^{k}Q_{0}(s_{i})-\frac{1}{n}\sum_{i=1}^{k}Q_{\theta}(s_{i})\right]\right\}.$
@@
8:Neural sub-networks $\{f_{\theta}^{(k)}\}_{k=0}^{K}$, pricing kernel function $Q_{\rho}(\cdot)$, benchmark function $Q_{0}(\cdot)$, utility function $U(\cdot)$, budget $\bar{x}$, sample size $n$, learning rate $\eta$, number of Adam steps $A_{s}$, penalty weights $w_{1},w_{2}$.
@@
  4. d
    In Steps 8-22, the approximation accuracy of the objective function and constraints can be improved by increasing the number of sampled points $n$. The Adam optimizer is used to update the network parameters, and the number of Adam steps $A_{s}$ determines how many gradient-based updates are performed during training, thus controlling the convergence of the network.
@@ 6.2 Experimental Results
To validate the model, we first conduct experiments under conditions in which the optimal solution $Q^{*}(\cdot)$ is available. The optimality is because Algorithm 1 results numerically satisfy the characterization in Theorem 2. We follow the setup in Section 5.4, which considers both the exponential utility and log utility under a linear SSD constraint quantile $Q_{0}$ (Table 8 (a)(c)).
@@
    Figure 6 (i)-(ii) Left: Compared with Figure 4 (a) and Figure 5 (d), the algorithm-guided piecewise-neural-network framework successfully preserves the structure features of the optimal solution, achieving a close match.
@@
    Figure 6 (i)-(ii) Right: The convergence speed depends on the number of sub-networks: the exponential case uses 2 sub-networks, while the log case uses 3, which also reflects the problem complexity. In simpler cases (Figure 6 (i)), the convergence curves are smooth and rapid because the two constraints in Problem (39) are quickly satisfied. In more complex cases (Figure 6 (ii)), the network spends more time mitigating constraint penalties, resulting in slower convergence.
Here we come to the conclusion that our approach, first deriving a sub-optimal solution via the proposed algorithm and then leveraging it to guide the neural network design, effectively captures the essential properties of the optimal solution and provides stable, high-quality performance for the SSD problem.
Extending to non-concave utility functions introduces a substantially more challenging problem. The difficulty stems from several factors. First, the optimization is infinite-dimensional over the space of admissible allocation functions, meaning that the solution is a functional rather than a finite-dimensional vector. Second, in the presence of SSD constraint and budget constraint, the feasible set becomes highly restricted and non-convex, which further complicates convergence.
Remarkably, our algorithm-guided piecewise-neural-network framework remains stable under these challenging conditions and achieves rapid convergence. By leveraging sub-optimal solutions obtained from our proposed Algorithm 1 as analytic priors and structuring the network in a piecewise manner, the framework effectively reduces the functional search space and guides the optimization toward regions that respect the SSD constraints. In contrast, a standard monolithic-neural-network (non-piecewise) framework exhibits substantially slower convergence, often requiring tens of times more training steps to reach convergence.
@@
Table 11: Convergence step and constraint satisfied step.
@@
    Table 11 (i) vs. (ii): The significant speedup under the piecewise framework primarily stems from rapid satisfaction of the penalty terms. By incorporating analytic prior inspired from Algorithm 1 into the network design, our piecewise framework enforces the SSD and budget constraints more efficiently, allowing the neural network to converge quickly while maintaining better approximation of the sub-optimal solution.
In summary, our study demonstrates the effectiveness of the algorithm-guided piecewise-neural-network framework for solving the SSD problem. For cases where an analytical optimal solution exists, our framework achieves results that closely match the optimal solution both numerically and structurally, indicating its high precision. From the perspective of convergence, our framework significantly accelerates the optimization process and is capable of escaping local optima, thereby attaining superior objective value compared with the standard monolithic-neural-network framework. These findings highlight the efficiency and robustness of our approach across different problem settings, and suggest strong potential for extension to more complex scenarios.

### v2

@@ 4.2 SSD Problem: Numerical Algorithm
    In Steps 8-12, we need to check whether the constructed $y_{\text{sub}}(\cdot)$ is non-decreasing on (0,1) to make the solution valid. It holds in many cases. In some invalid cases, we could still ensure the monotonicity by redesigning the construction of $t_{i}$ defined in Eq. (20), which is named $t_{\text{left}}$ in Eq. (23). The specific case will be shown in Section 5.
@@
    In Step 15, we need to check whether the condition holds numerically. It holds in many cases.
@@ 5 Numerical Results
Our study is motivated by the Black-Scholes model in a complete market. A classic Black-Scholes model consists of one riskless bond ($\frac{\mathrm{d}B_{t}}{B_{t}}=r\mathrm{d}t,~t\in[0,T]$, where the risk-free rate is $r=0.05$) and one stock ($\frac{\mathrm{d}S_{t}}{S_{t}}=\mu_{\text{S}}\mathrm{d}t+\sigma_{\text{S}}\mathrm{d}W_{t},~t\in[0,T]$, which represents a geometric Brownian motion with the expected return rate $\mu_{\text{S}}=0.086$, the volatility parameter $\sigma_{\text{S}}=0.3$ and a standard Brownian motion $\{W_{t}\}_{0\leqslant t\leqslant T}$). The wealth process is given by $\mathrm{d}X_{t}=\left(rX_{t}+(\mu_{S}-r)\pi_{t}\right)\mathrm{d}t+\sigma_{S}\pi_{t}\mathrm{d}W_{t},~t\in[0,T]$ and $X_{0}=\overline{x}$, where $\{\pi_{t}\}_{0\leqslant t\leqslant T}$ is the control process representing the investment amount in the stock and $T\in(0,\infty)$ is the evaluation time of investment. We define the pricing kernel process $\{\rho_{t}\}_{0\leqslant t\leqslant T}$ by
@@
As the market is complete and one can use the martingale method to duplicate the optimal portfolio process, it is sufficient to solve the optimal terminal wealth variable via the static problem (1) (see, e.g., Appendix A of Liang and Liu (2024)). Hence, our focus is solving the optimal wealth variable in Problem (1). Adapting to Problem (1), we denote the terminal variables of pricing kernel and wealth by $\rho:=\rho_{T}$ and $X:=X_{T}$, with a slight abuse of notation. Solving Eq. (29), $\rho$ follows the log-normal distribution (i.e., $\log(\rho)\sim\text{N}(-(r+\theta^{2}/2)T,(\theta\sqrt{T})^{2})$) and has a quantile function
$Q_{\rho}(t)=\exp\left\{\theta\sqrt{T}\Phi^{-1}(t)-(r+\theta^{2}/2)T)\right\}\triangleq\exp\left\{\sigma\Phi^{-1}(t)+\mu\right\},~t\in[0,1],$ (30)
@@ 5.1 FSD Problem: S-shaped Utility (Theorem 1)
We begin by specifying an S-shaped utility $U:[L,\infty)\rightarrow\mathbb{R}$, following the general S-shaped utility formulation in Definition 2:
@@
We then consider two portfolio selection problems. The first is Problem (3) with the general setting of S-shaped utility in Definition 2 (in particular, Eq. (31) with $L=$-$\infty$) and the benchmark quantile $Q_{0}(t)=10t^{2}-1,\;t\in[0,1].$ The second is the Merton problem (1) using the S-shaped utility setting in Eq. (31):
$\max_{X\in L^{0}}\mathbb{E}[U(X)],\text{ s.t. }\mathbb{E}[\rho X]\leqslant\overline{x},\;X\geqslant L\text{ a.s., }$ (32)
where the liquidation boundary is given by L=5L=-5 and $\bar{x}=5$. Here we add the liquidation boundary in order to make the second problem well defined and compare this classic solution with the first problem.
In this example, even though the first Problem (3) does not admit a liquidation constraint ($L=$-$\infty$), the first-order SD constraint acts a similar role: if $Q_{0}(1-F_{\rho}(\rho))=Q_{0}(0)$, then the optimal solution $X^{*}$ is located at the boundary $Q_{0}(1-F_{\rho}(\rho))$, otherwise $X^{*}$ is the same as the classic solution $(U_{1}^{\prime})^{-1}(\lambda\rho)$.
Figure 1: FSD problem: (31) vs (32).
@@ 5.2 SSD Problem: Power Utility
$Q_{0}(t)=\exp\left\{\sigma_{0}\Phi^{-1}(t)+\mu_{0}\right\},\;t\in[0,1].$ (34)
@@
Table 1: Parameter settings in the numerical illustration.
@@
Figure 2: SSD Problem: Setting (33) and (34).
@@
We apply the Poor-Performance-Region Algorithm (PPRA), and the explanations and financial insights from Figure 2 and Table 2 are as follows:
@@ 5.3 SSD Problem: S-shaped Utility
Table 3: Parameter settings in Figure 3.
@@
We apply the Poor-Performance-Region Algorithm, and the explanations and financial insights from Figure 3 and Table 4 are summarized as follows:
@@
    Figure 3: Across the S-shaped utility cases, we observe that when the market performs poorly, relative to the classical solution, the PPRA wealth exhibits a clear improvement, driven by the SSD constraint. Therefore, in an adverse market scenario, the SSD constraint effectively performs as a risk-control mechanism, ensuring that the PPRA wealth remains at least as high as the benchmark wealth.
@@
    Figure 3 (e) vs (f): In these examples, the poor performance region splits into two disjoint intervals. We further observe that when the budget of $Q_{0}$ is closer to the bound $\bar{x}$, the PPRA wealth becomes closer to the benchmark. This phenomenon arises because a high level of the budget of $Q_{0}$ makes the SSD constraint dominate the optimization.
@@ 5.4 SSD Problem: Various Utilities and Benchmark Quantiles
Based on the proposed algorithm, we further extend its applicability to a broader class of the SSD problems. To assess the generality of our approach, we conduct numerical experiments using different utilities and benchmark quantiles $Q_{0}$.
For the utility function $U(x)$, we consider several representative forms capturing different risk preferences, as shown in Table 5.
@@
Table 5: Settings of various utilities.
@@
Table 7: General settings in Figures 4-5.
@@
(e) piecewise log-normal 3.5 $\mu_{0}=-1,\sigma_{0}=3,k_{0}=2.3$ $p_{1}=q_{1}=0.6,p_{2}=0.8,$ $q_{2}=0.9,\lambda_{1}=1,\lambda_{2}=2$
(f) piecewise exponential 1.3 $\alpha=0.7,k_{0}=2.3$ $p_{1}=q_{1}=0.6,p_{2}=0.8,$ $q_{2}=0.9,\lambda_{1}=1,\lambda_{2}=2$
Table 8: Parameter settings in Figures 4-5.
Applying the Algorithm 1, we obtain the numerical results in Table 9 and plot Figures 4-5, where each panel illustrates the structure of the correction function $y_{\text{sub}}(\cdot)$ and the PPRA solution.
@@
Figure 5: Settings (d)(e)(f) in Table 9.
Later, we will clarify the structural behavior of the correction function $y_{\text{sub}}(\cdot)$. The numerical experiments yield the following explanations and financial insights:
@@
    Figure 4 (a)(b)(c): In a bad market scenario, the correction function $y_{\text{sub}}(\cdot)$ satisfies $y_{\text{sub}}(\cdot)=y_{0}(\cdot)$. Therefore, the PPRA wealth will coincide with the benchmark wealth. In this scenario, the benchmark serves as a risk-control mechanism, which is similar to Figure 2 (a)(b)(d).
@@
    Figure 4 (c)(f): Compared with Figure 4 (a)(b), the region where $y_{\text{sub}}(\cdot)=y_{0}(\cdot)$ expands. Hence, under relatively stagnant or bad market scenario, the PPRA wealth tends to coincide with the benchmark wealth. In such scenario, the SSD constraint plays a more prominent role by ensuring a benchmark wealth to reduce risk.
@@
    Figure 5 (d): The correction function $y_{\text{sub}}(\cdot)$ equals $y_{0}(\cdot)$ over certain intermediate intervals, causing the PPRA wealth to coincide with the benchmark in the mid-range market scenario, as shown in (d)’s right panel.
@@
    Figure 5 (e): The correction function $y_{\text{sub}}(\cdot)$ equals $y_{0}(\cdot)$ at both endpoints of the interval $(0,1)$. This implies the PPRA wealth must coincide with the benchmark wealth in both extremely bad and exceptionally favorable market scenarios.
In Figure 5 (e), we can observe how Steps 8-12 of Algorithm 1 ensure that the constructed function $y_{\text{sub}}(\cdot)$ remains non-decreasing on $(0,1)$. The key idea is to modify the construction of $g_{i}(t)$ in Eq. (19) so that both monotonicity requirements of $y_{\text{sub}}(\cdot)$ and the condition $z^{*}(\cdot)\leqslant 0$ are satisfied throughout the region in which $y_{\text{sub}}(\cdot)$ is updated. In this example, the interval where $y_{\text{sub}}(\cdot)=y_{\text{sub}}(t_{1})$ is extended into the region covered by the previous iteration (i.e., the first iteration), updating the original $y_{\text{sub}}(\cdot)$ in this region.
@@ 6 Algorithm-Guided Piecewise-Neural-Network Framework
In this section, we propose a novel approach to solve the SSD Problem (14) by designing a piecewise-neural-network-framework combined with Algorithm 1. The key observation is that in Algorithm 1, the poor performance region $C$ defined in Eq. (27) and the construction of the correction function $y_{\text{sub}}(\cdot)$ provide valuable structure information about the optimal solution, thereby guiding the design of an effective neural network framework.
@@ 6.1 Model Settings
where $f^{(k)}_{\theta_{k}}$ denotes the $k$-th neural sub-network parameterized by $\theta_{k}$, the parameters of the integrated network $Q$ are denoted as $\theta=\{\theta_{0},\theta_{1},\dots,\theta_{K}\}$ and $\mathbf{s}_{\text{feat}}(s)$ represents the Fourier feature embedding of the scalar input $s$ defined as
@@
$C_{1}=\frac{1}{n}\sum_{i=1}^{n}Q_{\theta}(s_{i})Q_{\rho}(1-s_{i}),\;\;C_{2}=\max\left\{0,\max_{k=1,\dots,n}\left[\frac{1}{k}\sum_{i=1}^{k}Q_{0}(s_{i})-\frac{1}{k}\sum_{i=1}^{k}Q_{\theta}(s_{i})\right]\right\}.$
@@
8:Neural sub-networks $\{f_{\theta_{k}}^{(k)}\}_{k=0}^{K}$, pricing kernel function $Q_{\rho}(\cdot)$, benchmark function $Q_{0}(\cdot)$, utility function $U(\cdot)$, budget $\bar{x}$, sample size $n$, learning rate $\eta$, number of Adam steps $A_{s}$, penalty weights $w_{1},w_{2}$.
@@
  4. d.
    In Steps 8-22, the approximation accuracy of the objective function and constraints can be improved by increasing the number of sampled points $n$. The Adam optimizer is used to update the network parameters, and the number of Adam steps $A_{s}$ determines how many gradient-based updates are performed during training, thus enhancing the convergence of the network.
@@ 6.2 Experimental Results
To validate the model, we first conduct experiments under conditions in which the optimal solution $Q^{*}(\cdot)$ is available. This optimality holds because Algorithm 1 results numerically follow the characterization in Theorem 1. We use the setup in Section 5.4, which considers both the exponential utility and log utility under a linear SSD constraint with benchmark quantile $Q_{0}$ (Table 8 (a)(c)).
@@
    Figure 6 (i)(ii)’s left panel: Compared with Figure 4 (a) and Figure 5 (d), the algorithm-guided piecewise-neural-network framework successfully preserves the structure features of the optimal solution, achieving a close match.
@@
    Figure 6 (i)(ii)’s right panel: The convergence speed depends on the number of sub-networks: the exponential case uses 2 sub-networks, while the log case uses 3, which also reflects the problem complexity. In simpler cases (Figure 6 (i)), the convergence curves are smooth and rapid because the two constraints in Problem (40) are quickly fulfilled. In more complex cases (Figure 6 (ii)), the network spends more time mitigating constraint penalties, resulting in slower convergence.
Here we come to the conclusion that our approach, first deriving a sub-optimal solution via the proposed algorithm and then leveraging it to guide the neural network design, effectively captures the essential properties of the optimal solution and provides stable, high-quality performance for the SSD problem with strictly with concave utilities.
Remarkably, our algorithm-guided piecewise-neural-network framework remains stable under a nonconcave utility and achieves rapid convergence in Figure 7. By leveraging sub-optimal solutions obtained from our proposed Algorithm 1 as analytic priors and structuring the network in a piecewise manner, the framework effectively reduces the functional search space and guides the optimization toward regions that fulfill the SSD constraints. In contrast, a standard monolithic-neural-network (non-piecewise) framework exhibits substantially slower convergence, often requiring tens of times more training steps to reach convergence.
@@
Table 11: Convergence steps and constraint satisfied steps.
@@
    Table 11 (i) vs (ii): The significant speedup under the piecewise framework primarily stems from rapid satisfaction of the penalty terms. By incorporating the analytic prior inspired from Algorithm 1 into the network design, our piecewise framework enforces the SSD and budget constraints more efficiently, allowing the neural network to converge quickly while maintaining better approximation of the sub-optimal solution.
In summary, our study demonstrates the effectiveness of the algorithm-guided piecewise-neural-network framework for solving the SSD problem. For cases where an analytical optimal solution exists, our framework achieves results that closely match the optimal solution both numerically and structurally, indicating its high precision. From the perspective of convergence, our framework significantly accelerates the optimization process and is capable of escaping local optima, thereby attaining superior performance compared with the standard monolithic-neural-network framework. Our findings highlight efficiency and robustness of our approach across different model settings and suggest strong potential for extension to more complex scenarios.